Please help. No book. This is a worksheet.
EVALUATE h(-2/3) if h(x)= [x-2]

Answer by Alan3354(21608)

EVALUATE h(-2/3) if h(x)= [x-2]
-------------------
Substitute -2/3 for x
h(-2/3) = -2/3 - 2
= -2 2/3
or -8/3